Find the largest integer k such that the equation
5x^2 - kx + 8 + 8x^2 - 17 = 0
has no real solutions.

Answer by ikleyn(52781)

.
Find the largest integer k such that the equation
5x^2 - kx + 8 + 8x^2 - 17 = 0
has no real solutions.
~~~~~~~~~~~~~~~~~~

Reduce to the standard form quadratic equation by combining like terms

    13x^2 - kx - 9 = 0.


Write the discriminant

    d = (-k)^2 - 4*13*(-9) = k^2 + 468.


Discriminant is always positive at any  real value of "k".
Therefore, given equation has real solution at any value of "k".


The given assignment is intently written in the form to confuse a reader.

Solved.